### v1

@@ 4.2 Regional Heterogeneity
Previous research has demonstrated that attitudes toward LGB people can substantially vary between countries, causing also a large variation in the number of individuals openly identifying with the LGB community [[, see, e.g.,]]badgett2020economic, badgett2021lgbtq.252525 An OECD report shows that even within a set of relatively comparable OECD countries, the size of the LBG communities differs by a factor of four [174]. To explore regional differences in players’ attitudes towards this community, we exploit the regional information of the matches. We divide our sample based on the server on which the matches were hosted. For comparability and data availability reasons, the matches are classified into four regional categories: European matches (North and East Europe and West Europe servers), Korean matches, Latin American matches (Brasil, North Latin America, and South Latin America servers), and North American matches. We then apply the synthetic control estimator of Section 3.2 to each of these categories separately.
Figure 4.2: Graves’ daily pick rates and synthetic control estimation results by region. The actual series are smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure 4.2 and Table 4.1 display the results.262626 Figure A.2 in Appendix A displays the identities and the contributions of the characters in the donor pool with non-zero estimated weights. The synthetic control estimator closely approximates the trajectory of Graves’ pick rates for matches in Europe and Latin America before the disclosure, exhibiting pre-treatment root mean squared errors comparable to those of the pooled specifications. However, discrepancies arise in Korean and North American matches, where the pre-treatment root mean squared error is two to three times higher than that achieved with European and Latin American matches. In Europe and Latin America, we estimate a negative and persistent effect of the coming-out event on players’ preferences for Graves. Point estimates are consistently negative across the considered specifications.
Although the magnitudes of the estimated effects vary across regions, it is important to consider that this does not necessarily reflect differential attitudes towards the LBG community, as various other factors may differ across servers. One such factor could be the differential levels of competitiveness on different servers, which may affect the character selection process by introducing different levels of subjectivity. We therefore carefully interpret our results in this section as evidence for anti-LGB sentiments in various regions, rather than interpreting the differences in the effect sizes.
@@ 5 Mechanisms
First, we examine the idea that shifts in character relative strengths could explain our estimated effect. We rule out this possibility in Section 5.1 by demonstrating that Graves’ strength remained unaffected by the coming-out event. Second, we explore the potential influence of players’ skills on their decision to abandon Graves. In Section 5.2, we show that players’ skills have no correlation with the choice to drop the character, thus dismissing the possibility that gameplay factors are the driving force behind the players’ observed behavior. Third, we investigate whether players transitioning away from Graves experience any performance-related consequences. This is the topic of Section 5.3, where we present evidence that switching to other characters does not affect the performance of the players involved. This emphasizes our ability to measure players’ true social attitudes and stigma, avoiding any potential biases stemming from strategic performance considerations.
Finally, we acknowledge that questions may arise about whether the findings of Section 4.1 are solely a consequence of Graves’ disclosure or if they are influenced by the broader context of LGBT Pride Month. We exploit the presence of other playable characters with sexual minority status to show that LGBT Pride Month is unlikely to explain our findings. Specifically, Appendix D introduces a theoretical framework that formalizes the existence of two “simultaneous treatments” and outlines sufficient assumptions that enable us to separate the impacts of coming out and LGBT Pride Month on players’ preferences for Graves. The results, detailed in Section 5.4, support the interpretation that the estimated effects are driven by Graves’ disclosure.
@@ 5.1 Graves’ Strength
To address this concern, we employ the synthetic control estimator described in Section 3.2 to examine the potential impact of the coming-out event on Graves’ strength. We measure characters’ strength using daily win rates, which indicate the percentage of matches won by a character out of the total matches they participated in each day.
Figure 5.1: Graves’ daily win rates and synthetic control estimation results. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure 5.1 displays the results. Overall, our analysis reveals that the coming-out event had no impact on Graves’ strength. Despite the actual series exhibiting daily fluctuations around the $50\%$ mark, the synthetic control estimator effectively captures its pre-treatment trend, showcasing its ability to predict the counterfactual trend. After the treatment date, the synthetic control estimator continues to align with Graves’ win rate trend, confirming that the character’s strength was unaffected by the disclosure. The average effect is estimated to be $-1.446$ percentage points (standard error: $3.506$), and the conventional $95\%$ confidence interval encompasses zero, indicating a failure to reject the null hypothesis of no effect. These findings demonstrate that Graves’ strength remained unchanged during the coming-out event, dismissing the possibility of a shift in his strength as an explanation for the results of Section 4.1.
Moreover, we note that players have real-time access to detailed information regarding characters’ strengths, weaknesses, and overall performance, as numerous websites continuously provide updated data on characters’ in-game statistics.272727 Examples of such websites include https://lolalytics.com/lol/graves/build/ and https://www.leagueofgraphs.com/champions/stats/graves. Therefore, players were well-informed that no game-relevant skills or attributes were altered during the treatment period, and they could observe that Graves’s strength remained consistent. These factors suggest that the negative impact of the coming-out event estimated in Section 4.1 is unlikely to be driven by actual or presumed changes in character relative strengths.
@@ 5.2 Players’ Skills
To address this concern, we examine the correlation between players’ skills and their decision to abandon Graves. We classify players into two groups based on their preferences for Graves before his disclosure: the first group comprises those who never selected Graves before the coming-out event (henceforth labeled as non-prior users), while the second group comprises those who chose Graves at least once in their matches before the coming-out event (henceforth labeled as prior users). We then examine performance differences both within and between these groups before and after the treatment. To mitigate potential noise from players with limited match appearances, we restrict our analysis to players who engaged in a minimum of $100$ matches before the disclosure. This yields a sample of $2872$ players, with $1834$ being non-prior users.
@@
Figure 5.2: Players’ average pick rates for Graves and performance measures. Players are divided into two groups based on their preferences for Graves before his disclosure. The panels display the average pick rates for Graves, number of daily matches, and win rate of each group before and after the coming-out event.
@@
The top panel of Figure 5.2 displays the average pick rate for Graves among prior and non-prior users before and after the treatment. We observe a sharp decline in pick rates among prior users following the coming-out event, similar in size to the decrease shown in Figure 4.1. Conversely, non-prior users exhibit a marginal increase in average pick rates post-treatment, although this increase is practically negligible.
@@
In the remaining panels of Figure 5.2, we investigate whether prior and non-prior users exhibit differences in their characteristics. First, the bottom left panel displays the average number of daily matches played by players. We observe similar numbers between groups both before and after the treatment, indicating that prior and non-prior users tend to engage in a comparable number of matches each day. We also note no differences in matches played post-treatment within both groups. This suggests that players are not leaving the game after the coming-out event. Instead, they are shifting their focus to other characters. Furthermore, Figure A.3 in Appendix A reveals that neither prior nor non-prior users change roles following Graves’ disclosure. This observation aligns with the intuition that players have sufficient alternatives, allowing them to easily switch to other characters and express their preferences at a low cost.
@@
Second, the bottom right panel displays the players’ average win rates, a metric capturing players’ skills by measuring the percentage of matches won out of their total engagements. We observe no substantial disparities within and between groups, indicating that the preference for Graves and the decision to abandon this character are unrelated to players’ skill levels. Overall, these findings dismiss the possibility that game-play factors are the driving force behind the estimated effects of Section 4.1, lending additional support to the social stigma attached to playing an LGB character as the mechanism underlying the players’ observed behavior.
@@ 5.3 Players’ Performance
We employ difference-in-differences identification and estimation strategies to assess the impact of players abandoning Graves on their performance. We gauge players’ performance by their daily win rate, which measures the percentage of matches won out of their total engagements. Our analysis focuses on the $1038$ prior-users of Section 5.2, who are classified into treated or control groups based on their responses to Graves’ disclosure. We consider different definitions of the treatment, sorted by their intensity. In the first version, labeled any reduction, we classify as treated those players who decreased their average pick rate for Graves following his disclosure, regardless of the extent (the number of treated units is $857$). In the second version, labeled substantial reduction, we classify as treated those players who reduced their average pick rate for Graves by at least $50\%$ post-disclosure (the number of treated units is $738$). In the third version, labeled complete abandonment, we classify as treated those players who exhibit a zero pick rate for Graves after the treatment (the number of treated units is $586$).
@@
Under the standard assumptions of parallel trends and no anticipation [[, see, e.g.,]]roth2023s, we can identify the average treatment effect on the treated (ATT) using observable data. The parallel trend assumption posits that the performance of treated and untreated players would have evolved similarly if Graves’ disclosure had not occurred. While we cannot formally test this assumption, the findings of Section 5.1 and Section 5.2 provide substantial support for its plausibility.282828 Moreover, we demonstrate below the absence of pre-treatment differences in trends by reporting placebo estimates of the ATT that are not statistically different from zero. This is often viewed as a natural plausibility check, although even if pre-trends are perfectly parallel, this does not necessarily guarantee the satisfaction of the post-treatment parallel trends assumption [[, see, e.g.,]]roth2023s. As for the no anticipation assumption, it stipulates that in the weeks preceding the disclosure, players’ performance did not change due to the incoming Graves’ disclosure. The plausibility of this assumption was thoroughly discussed in Section 2.2 and Section 4.1.
@@
We implement the approach of [186] to target the ATT at a particular day $t>T^{pre}$:292929 The framework outlined in [186] is broader as it accommodates multiple groups defined by the timing of treatment reception. This enables the identification and estimation of the group-time ATTs, defined as $ATT\left(g,t\right):=\operatorname{\mathbb{E}}\left[Y_{i,t}\left(g\right)-Y_{i,t}\left(0\right)|G_{g}=1\right]$, where $G_{g}$ is a binary variable indicating treatment reception in period $g$. However, our data set features a single group, given that all treated players receive the treatment simultaneously (i.e., at Graves’ disclosure date). This allows us to simplify notation and focus on the time ATTs in equation (5.1) for the single group we observe.
$ATT\left(t\right):=\operatorname{\mathbb{E}}\left[Y_{i,t}\left(1\right)-Y_{i,t}\left(0\right)|D_{i}=1\right]$ (5.1)
@@
where potential outcomes are defined as in Section 3.2, and $D_{i}$ is a binary variable indicating whether a player is treated or not. Under the assumptions of parallel trends and no anticipation, [186] show that $ATT\left(t\right)$ can be identified by comparing the change in outcomes between the latest period before the coming-out event and day $t$ experienced by treated players to the change in outcomes experienced by control players.303030 Formally, [186] show that $ATT\left(t\right)=\operatorname{\mathbb{E}}\left[Y_{i,t}-Y_{i,T^{pre}}|D_{i}=1\right]-\operatorname{\mathbb{E}}\left[Y_{i,t}-Y_{i,T^{pre}}|D_{i}=0\right]$. Estimation is carried out by replacing expectations with their sample analogs.
@@
Figure 5.3: Point estimates and simultaneous $95\%$ confidence bands allowing for clustering at the player level for the $ATT\left(t\right)$ estimated under an unconditional parallel trend (left panels) and a conditional parallel trend assumption (right panels). Red lines refer to pre-treatment periods, while blue lines refer to post-treatment periods. Each row corresponds to a different version of the treatment.
@@
The left panels of Figure 5.3 display the point estimates and simultaneous $95\%$ confidence bands for the $ATT\left(t\right)$.313131 Table A.1 in Appendix A reports aggregated results by displaying the average time ATT. Overall, we find that shifting away from Graves to other characters has no impact on players’ performance. None of the estimated $ATT\left(t\right)$ is statistically different from zero, suggesting that transitioning to other characters does not result in any performance-related consequences. This finding highlights that the decision to move away from Graves is not influenced by performance considerations.
As a robustness check, we explore an alternative scenario where the parallel trends assumption is required to hold only conditional on pre-treatment covariates. In this context, we identify and estimate $ATT\left(t\right)$ using the doubly-robust approach of [186].323232 In essence, this approach entails estimating the change in outcomes for control players conditional on the pre-treatment covariates $X_{i}$ and averaging out $X_{i}$ over the distribution of covariates for treated players. For a more detailed understanding of this approach, readers are referred to [186]. Our pre-treatment covariates encompass players’ skills information, such as average kills, deaths, assists, and gold earned prior to the treatment, as well as the average number of daily matches they engaged in before the treatment. The right panels of Figure 5.3 display the results. The results are consistent with those obtained under the unconditional parallel trend assumption.
@@
Finally, Figure 5.3 also displays placebo estimates of the time ATTs for the ten days before the treatment.333333 Figure A.4 in Appendix A displays the remaining estimated placebo $ATT\left(t\right)$. As explained above, these estimates are valuable for “pre-testing” the credibility of the parallel trend assumption [186]. Notably, all placebo time ATTs in the pre-treatment periods are statistically insignificant, supporting the validity of the parallel trends assumption.

### v3

@@ 4.3 Regional Heterogeneity
Previous research has demonstrated that attitudes toward LGB people can substantially vary between countries, causing also a large variation in the number of individuals openly identifying with the LGB community [[, see, e.g.,]]badgett2020economic, badgett2021lgbtq.313131 An OECD report shows that even within a set of relatively comparable OECD countries, the size of the LBG communities differs by a factor of four [171]. To explore regional differences in players’ attitudes towards this community, we exploit the geographic information associated with each match. Specifically, we divide the sample according to the macro-regions defined in Section 3.1 and apply the synthetic control estimator described in Section 3.2 separately within each region.
Figure 4: Graves’ daily pick rates and synthetic control estimation results by region. The actual series are smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure 4 displays the results.323232 Figure A.V in Appendix A displays the identities and the contributions of the characters in the donor pool with non-zero estimated weights. The synthetic control estimator closely approximates the trajectory of Graves’ pick rates for matches in Europe and Latin America before the disclosure, exhibiting pre-treatment root mean squared errors comparable to those of the pooled specifications. However, discrepancies arise in Korean and North American matches, where the pre-treatment root mean squared error is two to three times higher than that achieved with European and Latin American matches. In Europe, Korea, and Latin America, we estimate negative and persistent effects of the coming-out event on players’ preferences for Graves. Point estimates for the average treatment effect in these regions, reported in Table A.II in Appendix A, are consistently negative and are statistically significant at the 5% level across most of the considered specifications. In North America, where the pre-treatment root mean squared is relatively large, we are unable to determine the sign of the impact, as the confidence intervals for the estimated average effect consistently encompass zero.
The estimated average effect varies substantially across regions, with the largest effect observed in Europe ($62.23\%$ relative to the pre-treatment average preferences) and the smallest effect observed in Korea ($31.82\%$) in our main specification, indicating heterogeneous responses to the coming-out event. However, it is important to consider that the different magnitudes of the estimated effects do not necessarily reflect differential attitudes towards the LBG community, as various other factors may differ across servers. One such factor could be the differential levels of competitiveness on different servers, which may affect the character selection process by introducing different levels of subjectivity. We therefore carefully interpret our results in this section as evidence for anti-LGB sentiments in various regions, rather than interpreting the differences in the effect sizes.
@@ 5 Mechanisms
First, we examine the idea that shifts in character relative strengths could explain our estimated effect. We rule out this possibility in Section 5.1 by demonstrating that Graves’ strength remained unaffected by the coming-out event. Second, we explore the potential influence of players’ skills on their decision to abandon Graves. In Section 5.2, we show that players’ skills have no correlation with the choice to drop the character, thus dismissing the possibility that gameplay factors are the driving force behind the players’ observed behavior. Third, we investigate whether players transitioning away from Graves experience any performance-related consequences. This is the topic of Section 5.3, where we present evidence that switching to other characters does not affect the performance of the players involved. This emphasizes our ability to measure players’ true social attitudes and stigma, avoiding any potential biases stemming from strategic performance considerations. Fourth, in Section 5.4, we dismiss the possibility that the release of a new potentially substitute character drives the players’ decisions to switch away from Graves.
Finally, we acknowledge that questions may arise about whether the findings of Section 4 are solely a consequence of Graves’ disclosure or if they are influenced by the broader context of LGBT Pride Month. We exploit the presence of other playable characters with sexual minority status to show that LGBT Pride Month is unlikely to explain our findings. Specifically, Appendix D introduces a theoretical framework that formalizes the existence of two “simultaneous treatments” and outlines sufficient assumptions that enable us to separate the impacts of coming out and LGBT Pride Month on players’ preferences for Graves. The results, detailed in Section 5.5, support the interpretation that the estimated effects are driven by Graves’ disclosure.
@@ 5.1 Graves’ Strength
To address this concern, we employ the synthetic control estimator described in Section 3.2 to examine the potential impact of the coming-out event on Graves’ strength. We measure characters’ strength using daily win rates.
Figure 5: Graves’ daily win rates and synthetic control estimation results. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure 5 displays the results. Overall, our analysis reveals that the coming-out event had no impact on Graves’ strength. Despite the actual series exhibiting daily fluctuations around the $50\%$ mark, the synthetic control estimator effectively captures its pre-treatment trend, showcasing its ability to predict the counterfactual trend. After the treatment date, the synthetic control estimator continues to align with Graves’ win rate trend, confirming that the character’s strength was unaffected by the disclosure. The average effect is estimated to be 1.046-1.046 percentage points (standard error: $5.298$), and the conventional $95\%$ confidence interval encompasses zero, indicating a failure to reject the null hypothesis of no effect. These findings demonstrate that Graves’ strength remained unchanged during the coming-out event, dismissing the possibility of a shift in his strength as an explanation for the results of Section 4.
Moreover, we note that players have real-time access to detailed information regarding characters’ strengths, weaknesses, and overall performance, as numerous websites continuously provide updated data on characters’ in-game statistics.333333 Examples of such websites include https://lolalytics.com/lol/graves/build/ and https://www.leagueofgraphs.com/champions/stats/graves. Therefore, players were well-informed that no game-relevant skills or attributes were altered during the treatment period, and they could observe that Graves’s strength remained consistent. These factors suggest that the negative impact of the coming-out event estimated in Section 4 is unlikely to be driven by actual or presumed changes in character relative strengths.
@@ 5.2 Players’ Skills
To address this concern, we examine the correlation between players’ skills and their decision to abandon Graves. We classify players into two groups based on their preferences for Graves before his disclosure: the first group comprises those who chose Graves in at least $5\%$ of their matches before the coming-out event (henceforth labeled as prior users, $n=953$), while the second group comprises the remaining players (henceforth labeled as non-prior users, $n=6468$). We then examine performance differences both within and between these groups before and after the treatment.
@@
Figure 6: Players’ average pick rates for Graves and performance measures. Players are divided into two groups based on their preferences for Graves before his disclosure. The panels display the average pick rates for Graves, number of daily matches, and win rate of each group before and after the coming-out event.
@@
The top panel of Figure 6 displays the average pick rate for Graves among prior and non-prior users before and after the treatment. We observe a sharp decline in pick rates among prior users following the coming-out event, similar in size to the decrease shown in Figure 2 and Table 1. Conversely, non-prior users exhibit a marginal increase in average pick rates post-treatment, although this increase is practically negligible.
@@
In the remaining panels of Figure 6, we investigate whether prior and non-prior users differ in their in-game behavior. First, the bottom left panel shows the average number of daily matches played by each group. We observe similar engagement levels across groups, both before and after the treatment, indicating that prior and non-prior users tend to play a comparable number of matches per day. Notably, neither group appears to disengage from the game following Graves’ disclosure. On the contrary, the number of daily matches slightly increases post-treatment for both groups, suggesting that players are not leaving the game but instead redirecting their attention to other characters. Supporting this interpretation, Figure A.VI in Appendix A shows no meaningful change in role preferences for either group following the event. This pattern is consistent with the idea that players face low switching costs and have access to suitable alternatives, enabling them to express their preferences without disrupting their gameplay. These results align with those displayed in Table 1.
@@
Second, the bottom right panel displays the players’ average win rates. We observe no substantial disparities within and between groups, indicating that the preference for Graves and the decision to abandon this character are unrelated to players’ skill levels. Overall, these findings dismiss the possibility that game-play factors are the driving force behind the estimated effects of Section 4, lending additional support to the social stigma attached to playing an LGB character as the mechanism underlying the players’ observed behavior.
@@ 5.3 Players’ Performance
We employ difference-in-differences identification and estimation strategies to assess the impact of players abandoning Graves on their performance. We gauge players’ performance by their daily win rate. Our analysis focuses on the $953$ prior-users of Section 5.2, who are classified into treated or control groups based on their responses to Graves’ disclosure. We consider two different definitions of the treatment, sorted by their intensity. In the first version, labeled moderate reduction, we classify as treated those players who decreased their average pick rate for Graves following his disclosure by at most $75\%$ of their pre-treatment average pick rate (the number of treated units is $287$). In the second version, labeled substantial reduction, we classify as treated those players who reduced their average pick rate for Graves by $75\%$ to $100\%$ post-disclosure (the number of treated units is 491491).343434 Figure A.VII in Appendix A shows the distribution of percentage reduction in average pick rates for Graves from pre-treatment to post-treatment among prior users. In both scenarios, the control group remains the same and consists of the $175$ prior users who did not reduce their average pick rate for Graves following his disclosure.
@@
Under the standard assumptions of parallel trends and no anticipation [[, see, e.g.,]]roth2023s, we can identify the average treatment effect on the treated (ATT) using observable data. The parallel trend assumption posits that the performance of treated and untreated players would have evolved similarly if Graves’ disclosure had not occurred. While we cannot formally test this assumption, the findings of Section 5.1 and Section 5.2 provide substantial support for its plausibility.353535 Moreover, we demonstrate below the absence of pre-treatment differences in trends by reporting placebo estimates of the ATT that are not statistically different from zero. This is often viewed as a natural plausibility check, although even if pre-trends are perfectly parallel, this does not necessarily guarantee the satisfaction of the post-treatment parallel trends assumption [[, see, e.g.,]]roth2023s. As for the no anticipation assumption, it stipulates that in the weeks preceding the disclosure, players’ performance did not change due to the incoming Graves’ disclosure. The plausibility of this assumption was thoroughly discussed in Section 2.2 and Section 4.1.
@@
We implement the approach of [181] to target the ATT at a particular day $t>T^{pre}$:363636 The framework outlined in [181] is broader as it accommodates multiple groups defined by the timing of treatment reception. This enables the identification and estimation of the group-time ATTs, defined as $ATT\left(g,t\right):=\operatorname{\mathbb{E}}\left[Y_{i,t}\left(g\right)-Y_{i,t}\left(0\right)|G_{g}=1\right]$, where $G_{g}$ is a binary variable indicating treatment reception in period $g$. However, our data set features a single group, given that all treated players receive the treatment simultaneously (i.e., at Graves’ disclosure date). This allows us to simplify notation and focus on the time ATTs in equation (6) for the single group we observe.
$ATT(t):=\operatorname{\mathbb{E}}[Y_{i,t}\left(1\right)-Y_{i,t}\left(0\right)|D_{i}=1],$ (6)
@@
where potential outcomes are defined as in Section 3.2, and $D_{i}$ is a binary variable indicating whether a player is treated or not. Under the assumptions of parallel trends and no anticipation, [181] show that $ATT(t)$ can be identified by comparing the change in outcomes between the latest period before the coming-out event and day $t$ experienced by treated players to the change in outcomes experienced by control players.373737 Formally, [181] show that $ATT(t)=\operatorname{\mathbb{E}}[Y_{i,t}-Y_{i,T^{pre}}|D_{i}=1]-\operatorname{\mathbb{E}}[Y_{i,t}-Y_{i,T^{pre}}|D_{i}=0]$. Estimation is carried out by replacing expectations with their sample analogs.
@@
Figure 7: Point estimates and simultaneous $95\%$ confidence bands allowing for clustering at the player level for the $ATT(t)$. Red lines refer to pre-treatment periods, while blue lines refer to post-treatment periods. Each row corresponds to a different version of the treatment.
@@
Figure 7 displays the point estimates and simultaneous $95\%$ confidence bands for the $ATT(t)$. Overall, we find that shifting away from Graves to other characters has no impact on players’ performance. None of the estimated $ATT(t)$ is statistically different from zero, suggesting that transitioning to other characters does not result in any performance-related consequences. This finding highlights that the decision to move away from Graves is not influenced by performance considerations.
@@
Figure 7 further displays placebo estimates of the time ATTs for the ten days before the treatment.383838 Figure A.VIII in Appendix A displays the remaining estimated placebo $ATT(t)$. As explained above, these estimates are valuable for “pre-testing” the credibility of the parallel trend assumption [181]. Notably, all placebo time ATTs in the pre-treatment periods are statistically insignificant, supporting the validity of the parallel trends assumption.
5.4 New Substitute Character
On June $9^{th}$, $2022$, a new character (Bel’Veth) was released.393939 Figure A.IX in Appendix A displays Bel’Veth’s daily pick rates. Since the primary position the character is designed for is the same position as Graves is designed for, it can be considered a close substitute. Therefore, players’ decisions to switch away from Graves (after June $9^{th}$) might to some degree be driven by the desire to experiment with the new character and to explore potential competitive advantages, challenging social stigma as the primary explanation behind our main result.
Figure 8: Average pick rate for Bel’Veth after its release among prior users.
If the release serves as the primary explanation for the observed drop in Graves pick rate we would expect a positive correlation between the size of the drop in players’ Graves usage and players preferences for the new character. In Figure 8 we show that this is not the case. Figure 8 shows the average pick rate for Bel’Veth for the prior users of Section 5.2, classified as in Section 5.3. We find that those players who feature a substantial reduction in their Graves’ pick rates following the coming-out event are less likely to pick Bel’Veth for their matches than players who feature a moderate reduction. Moreover, these players are even less likely to pick Bel’Veth for their matches than players who did not react at all to Graves’ disclosure. We therefore argue that the release is unlikely to serve as the primary explanation for the observed main result of the paper.